OpenStudy (anonymous):

there are n teams in a league and the number of games played is given by the polynomial n2 - n = N. a woman's softball league has 11 teams what is the total number of games to be played?

OpenStudy (dumbcow):

11^2 - 11 = 121 - 11 = 110

OpenStudy (anonymous):

dumbcow help pls factor b3 + 27

OpenStudy (dumbcow):

use general formula: (a^3 + b^3) = (a+b)(a^2 - ab + b^2) here a = b. b = 3

OpenStudy (anonymous):

dumbcow this is the ans. i came up with b3 + 27 = (b+3)(b2-3b+9)

OpenStudy (dumbcow):

perfect

OpenStudy (anonymous):

thnks